help me solve x, if x(x-1)=0
Answer by Theo(13342)

your formula is x * (x-1) = 0

this equation will be true if x = 0 or if x = 1.

if x = 0, you will get 0 * -1 = 0

if x = 1, you will get 1 * 0 = 0

anything times 0 will always be equal to 0.

if you graph this equation by setting y = x * (x-1), your graph will cross the x-axis at the points 0 and 1.

this means that y will be equal to 0 at those values of x.

here's your graph.

$$$